OpenStudy (anonymous):

What would the unit rates be for a temperature change of 20 degrees in 4 hours? Record any fractional answers in decimal form @igreen

OpenStudy (anonymous):

plzz help @iGreen

OpenStudy (anonymous):

@Hero

OpenStudy (mathstudent55):

Unit rate means per one unit. In this case, it means degrees per hour. "per" means division, so you need to divide the number of degrees by the number of hours.

OpenStudy (mathstudent55):

In other words, 20 degrees in 4 hours is how many degrees in 1 hour?

OpenStudy (anonymous):

idk im really bad at math @mathstudent55

OpenStudy (mathstudent55):

Do you have difficulty understanding this statement: "divide the number of degrees by the number of hours." ?

OpenStudy (anonymous):

Is it 5 degrees/hours and 0.4 for hours/degree?

OpenStudy (anonymous):

if i gave you 20 sweets in 4 hours how many would you get per hour ?

OpenStudy (anonymous):

5?@jamaica25

OpenStudy (anonymous):

@jamaica25

OpenStudy (anonymous):

yes same as your question 5 deg.per hour

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

\(\dfrac{20^o}{4~hour} = 5~\dfrac{^o}{hour}\) \(\dfrac{4~hour}{20^o} = 0.2~ \dfrac{hour}{^o}\)